An acid, HAc, has pKa = 0. To a solution with a total volume of 0.50 dm. 0.500 moles of the acid and 0.500 moles of the sodium salt of the acid are added which is very easily soluble. What is the pH of the solution?

The answer is 0,38 or 0,4

Please help :(
Answer #1

pKa of HA = 0.0 According Henderson-Hasselbalch equation salt] pH = pka+log acid pH = pka + log [HA] A- =pKa + log [HA] =00+